# Understanding the Limits of Lifelong Knowledge Editing in LLMs

### v1

@@ Abstract
Keeping large language models factually up-to-date is crucial for deployment, yet costly retraining remains a challenge. Knowledge editing offers a promising alternative, but methods are only tested on small-scale or synthetic edit benchmarks. In this work, we aim to bridge research into lifelong knowledge editing to real-world edits at practically relevant scale. We first introduce WikiBigEdit; a large-scale benchmark of real-world Wikidata edits, built to automatically extend lifelong for future-proof benchmarking. In its first instance, it includes over 500K question-answer pairs for knowledge editing alongside a comprehensive evaluation pipeline. Finally, we use WikiBigEdit to study existing knowledge editing techniques’ ability to incorporate large volumes of real-world facts and contrast their capabilities to generic modification techniques such as retrieval augmentation and continual finetuning to acquire a complete picture of the practical extent of current lifelong knowledge editing.
@@ 1 Introduction
Keeping large language models factually up-to-date is a critical challenge in their practical deployment (Hartvigsen et al., 2023; Meng et al., 2023, 2022; Wang et al., 2024a; Zhang et al., 2024) due to given knowledge cutoffs and limited tail-end concept coverage (Cheng et al., 2024; Zheng et al., 2024; Zhang et al., 2024; Udandarao et al., 2024) incurring inaccuracies. Finding ways to inject or update new and revised facts is crucial; Particularly for factuality-critical applications like medicine (Bao et al., 2023; Jeong et al., 2024), law (Colombo et al., 2024; Lai et al., 2024) or education (Team et al., 2024c; Wang et al., 2024b). However, frequent full model retraining is prohibitively expensive. To address this, lifelong knowledge editing techniques (Hartvigsen et al., 2023; Meng et al., 2023; Wang et al., 2024a) have recently emerged. In controlled benchmark studies, they have shown success incorporating new facts along three key metrics: (1) generalization beyond merely memorizing question-answer pairs, (2) locality of edits to maintain pretraining knowledge, and (3) retention of all previously seen edits. However, existing benchmarks are small [^1] CounterFact (Meng et al., 2023) at 20K, ZsRE (Levy et al., 2017) at 1K, SelfCheckGPT (Manakul et al., 2023) at 600., synthetic, and precede the knowledge cutoff of modern LLMs. While more recent efforts Jang et al. (2022); Khodja et al. (2024) avoid synthetic edits by using wikidata, they remain small-scale (around 20K), and have yet to examine difficulty of edits, limiting our understanding of really lifelong knowledge editing capabilities at practically-needed scales. LLMs are nowadays trained on trillions of text tokens (e.g. Llama-3 (Grattafiori et al., 2024) on 15T), and studying lifelong factuality updates demands large volumes of recent real-world knowledge edits. To better understand lifelong knowledge editing at a scale and scenario of practical importance, we introduce WikiBigEdit - a large-scale, lifelong benchmark (Prabhu et al., 2024; Ghosh et al., 2024) with a fully automated dataset extraction pipeline that continuously expands with new real-world factual edits to ensure future-proof factuality probing. At its inception, WikiBigEdit covers over 500K high-quality question-answer pairs (around 7 million tokens; magnitudes larger than standard edit benchmarks) spanning eight time-intervals over five months (February - July 2024) - reflecting a typical interval between subsequent version releases of large-scale LLMs [^2] E.g. Llama 1 (Touvron et al., 2023a) to 2 (Touvron et al., 2023b), DeepSeek-1 (DeepSeek-AI et al., 2024a) to V2 (DeepSeek-AI et al., 2024b) and V3 (DeepSeek-AI et al., 2024c). to better emulate the deployment time required to cover with factual updates. Derived from periodic changes to Wikidata knowledge graphs (Jang et al., 2022; Khodja et al., 2024), WikiBigEdit covers a large range of factual edits and refinements. Moreover, WikiBigEdit introduces comprehensive evaluation axes going beyond standard knowledge editing performance measures, incorporating locality checks, multi-hop reasoning across incorporated edits as well as standard and complex generalization tests. We use WikiBigEdit to explore the properties of real-world knowledge edits, and conduct an extensive study of knowledge editing techniques for WikiBigEdit-scale updates, focusing on their ability to integrate numerous sequential edits while maintaining previous knowledge and factual accuracy beyond simple memorization. Additionally, we compare these methods against other generic modification techniques such as retrieval augmentation (Vu et al., 2023; Prabhu et al., 2023a; Gui et al., 2024) and continual finetuning (Ibrahim et al., 2024; Roth et al., 2024). Our study reveals limited transfer of knowledge editing techniques to practically scalable lifelong updates and highlights that even when evaluated around knowledge editing desiderata, standard approaches such as retrieval augmentation or continual finetuning with model merging can perform better.
@@
(1) We present WikiBigEdit: a large-scale benchmark of 500k high-quality question-answer pairs spanning eight timesteps over six months, derived from real-world factual changes in Wikidata to both analyze and simulate real-world knowledge edit tasks at a practically relevant scale.
(2) A fully automated extraction pipeline which continuously extracts new, suitable factual edits from Wikidata knowledge graphs, enabling the benchmark to evolve with new knowledge updates continuously.
(3) Using WikiBigEdit, we thoroughly analyze the capability of existing lifelong knowledge editing methods to conduct lifelong edits at scale; contrasted against retrieval augmentation and continual finetuning to understand limits in relation to other established approaches. Code at http://anonymous.4open.science/r/WikiBigEdit-70F1.
@@ 2 Related Works
NQ 307K 2016 Google Search queries Open-domain factual QA
@@
EX-FEVER 60K 2023 Hyperlinked Wikipedia Multi-hop fact verification
ELI5 270K 2019 Reddit (ELI5 subreddit) Long-form explanatory QA
@@
DatedData 200K Various Varied web sources Temporal QA (time-sensitive)
StreamingQA 150K 2007-2020 WMT news articles Real-time event-based QA
@@
CLARK 1.4K 2021-2024 Wikipedia Knowledge-intensive QA
@@
TemporalWiki 7K 2021 Wikipedia, Wikidata Temporal QA (time-sensitive)
@@
WikiBigEdit 500K+ 2024 Wikidata Fact-based QA
@@
(Lifelong) Knowledge Editing Benchmarks. Traditional QA datasets like NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and MS MARCO (Bajaj et al., 2018) target open-domain factual QA, HotpotQA (Yang et al., 2018) and EX-FEVER (Ma et al., 2024) assess multi-hop reasoning. Fan et al. (2019); Li et al. (2024); Cheng et al. (2024); Mallen et al. (2023) focus on long-form explanatory questions and time-sensitive QA. StreamingQA (Liška et al., 2022), ArchivalQA (Wang et al., 2021) or HelloFresh (Franzmeyer et al., 2024) address evolving or time-sensitive, event-driven data. Knowledge editing benchmarks assess the ability to integrate new facts without retraining. ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023) focus on synthetic updates, while AToKE (Yin et al., 2024) and ChronoEdit (Ge et al., 2024) evaluate temporal consistency but remain limited in scale and scope. ELKEN (Peng et al., 2024) and EVEDIT (Liu et al., 2024b) generate edits based on real-world events. TemporalWiki (Jang et al., 2022) and WikiFactDiff (Khodja et al., 2024), are more closely aligned with WikiBigEdit. Both evaluate temporal update consistency using consecutive Wikidata snapshots. WikiBigEdit provides over a magnitude more samples (Table 2), future-proofing and a comprehensive evaluation along all editing desiderata.
@@
Knowledge Editing. Knowledge editing methods include local modification, such as ROME (Meng et al., 2023) and MEMIT (Meng et al., 2022), Global Optimization like Mitchell et al. (2021), and External Memorization such as Mitchell et al. (2022); Hartvigsen et al. (2023); Chen et al. (2024) or WISE (Wang et al., 2024a). While these methods enable efficient updates, they face challenges like catastrophic forgetting and knowledge distortion during larger-scale and sequential edits (Hsueh et al., 2024; Gupta et al., 2024b; Yang et al., 2024b). Our benchmark provides a dynamic research framework to study this in-depth under real-world conditions.
Retrieval Augmented Generation (RAG) enhances pre-trained generative models with retrieval to address outdated knowledge and traceability (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020; Rashkin et al., 2022; Honovich et al., 2022; Borgeaud et al., 2022; Izacard et al., 2022). RAG also enables dynamic updates, reducing context-memory conflicts when parametric knowledge becomes outdated (Vu et al., 2023). However, RAG faces limitations in multi-hop reasoning and inference costs.
@@
Continual Finetuning and Model Merging. Continual finetuning has emerged as a reliable approach for general large-scale model updates (Garg et al., 2024; Prabhu et al., 2023b; Ibrahim et al., 2024). Roth et al. (2024) show that parameter-efficient finetuning via e.g. LoRA (Hu et al., 2021) or GaLore (Zhao et al., 2024), can be very effective, leveraging regularization provided by limiting trainable parameter counts (Thede et al., 2024). Model merging (Wortsman et al., 2022; Ilharco et al., 2022; Ramé et al., 2024) has shown promising gains coupled with finetuning for continual tasks (Stojanovski et al., 2022; Marouf et al., 2024; Kozal et al., 2024; Marczak et al., 2024; Dziadzio et al., 2024). We investigate and contrast these approaches against knowledge editing for large-scale knowledge editing.
@@ 3 WikiBigEdit
WikiBigEdit follows a three-pronged agenda: (1) extract a large number of realistic knowledge edits, (2) incorporate extensive evaluations to study generalization, reasoning and locality of edits, and (3) do so fully automated. The full pipeline is visualized in Figure 1.
Lifelong Knowledge Editing - Notation.
The task of large-scale, lifelong knowledge editing requires updating a pre-trained LLM $f^{0}$ with a stream of factual updates $[u_{1},u_{2},\ldots,u_{T}]$, where each update $u_{t}=(q_{t},a_{t})$ consists of a question $q_{t}$ and its factual answer $a_{t}$. These updates are grouped into $B$ sequential batches $[U_{1},U_{2},\ldots,U_{B}]$, where each batch can encompass anything from a single edit to multiple. For each batch update $b$ (also denoted as timestep in this work), the model $f^{b-1}$, trained on updates from prior batches $U_{<b}$, is further updated with the current batch $U_{b}$ to produce $f^{b}$. The model must meet three key objectives: integrate updates from $U_{b}$, retain knowledge from $U_{<b}$, and ensure that semantic understanding required to correctly solve previously seen edits generalizes beyond simple memorization of exact question-answer pairings.
@@ 3.1 Lifelong Automated Extraction of Factual Edits
As a proxy for world knowledge, we follow Jang et al. (2022); Khodja et al. (2024) and leverage Wikidata knowledge graphs. By performing pairwise comparisons of snapshots taken at different points in time (Jang et al., 2022), we identify changes that approximate real-world knowledge evolution. Given this, the automated generation of WikiBigEdit leverages the following pipeline:
@@
(5) Generation of Question-Answer Edit Pairs. For $\mathcal{S}_{\text{changed}}$, $\mathcal{S}_{\text{locality}}$ and $\mathcal{S}_{\text{mhop}}$, we generate corresponding question-answer pairs using GPT-3.5 (Brown et al., 2020) (GPT-4o-mini for multi-hop) through few-shot prompting (for all prompt templates please see the supplementary). For $\mathcal{S}_{\text{mhop}}$, we generate questions that omit the middle entity of the two-hop factual quintuple. Using this, we generate the corresponding question-answer sets $\mathcal{S}^{\text{QA}}_{\text{changed}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$. $\mathcal{S}^{\text{QA}}_{\text{locality}}$ and $\mathcal{S}^{\text{QA}}_{\text{mhop}}$ are used for evaluation, while $\mathcal{S}^{\text{QA}}_{\text{changed}}$ constitutes the respective fact-based training data.
@@ 3.2 Analysis and Exploration
In this section, we first summarize various benchmark characteristics, before leveraging the first instantiation of our large-scale WikiBigEdit to study the structure and properties of real-world knowledge edits derived from Wikidata knowledge graphs. Table 2 provides an overview of the eight time-intervals that constitute WikiBigEdit. The factual updates are extracted from Wikidata snapshots spanning February 2024 to July 2024 (extracted on a roughly bi-monthly frequency); with the availability of Wikidata snapshots determining size and extend of factual updates associated with each timestep. Filtered updates range from 27k to 69k per step, with monthly factual changes exceeding 100k. This variability reflects the dynamic nature of knowledge updates and practical constraints when aiming to ensure that models remain up-to-date. Overall, this version contains 506,035 question-answer pairs across 6.9M tokens.
@@
Specificity of factual updates. We analyze the specificity of knowledge edits, distinguishing between well-known and highly specialized facts. We use two measures: (1) the Infinigram API (Liu et al., 2024a) to estimate entity frequency in a proxy LLM training corpus, and (2) Wikipedia view counts from Jan 2023 to June 2024 (Mallen et al., 2023). As shown in Figure 2 (d-g), WikiBigEdit contains both general questions (e.g., Which continent is Krasnodar a part of?) and highly specific ones (e.g., Who was the doctoral advisor of Omar Khayyám?). To assess specificity’s impact on performance, we evaluate five LLMs (Llama-2-7b, Llama-3-8b, Mistral-7b, Gemma-7b, and xGen-7b) and correlate their accuracy. As shown in (h), lower corpus frequency (higher specificity) correlates with reduced performance, in line with Udandarao et al. (2024); Cheng et al. (2024). However, Wikipedia view counts (i) show inconsistent correlations. Finally, we note that question specificity appears primarily determined by the object rather than the subject. See supplementary for a more detailed analysis.

### v2

# WikiBigEdit: Understanding the Limits of Lifelong Knowledge Editing in LLMs

@@ Abstract
Keeping large language models factually up-to-date is crucial for deployment, yet costly retraining remains a challenge. Knowledge editing offers a promising alternative, but methods are only tested on small-scale or synthetic edit benchmarks. In this work, we aim to bridge research into lifelong knowledge editing to real-world edits at a practically relevant scale. We first introduce WikiBigEdit; a large-scale benchmark of real-world Wikidata edits, built to automatically extend lifelong for future-proof benchmarking. In its first instance, it includes over 500K question-answer pairs for knowledge editing alongside a comprehensive evaluation pipeline. Finally, we use WikiBigEdit to study existing knowledge editing techniques’ ability to incorporate large volumes of real-world facts and contrast their capabilities to generic modification techniques such as retrieval augmentation and continual finetuning to acquire a complete picture of the practical extent of current lifelong knowledge editing. [^1] Code available at https://github.com/ExplainableML/WikiBigEdit.
@@ 1 Introduction
Keeping large language models factually up-to-date is a critical challenge in their practical deployment (Hartvigsen et al., 2023; Meng et al., 2023, 2022; Wang et al., 2024a; Zhang et al., 2024) due to given knowledge cutoffs and limited tail-end concept coverage (Cheng et al., 2024; Zheng et al., 2024; Zhang et al., 2024; Udandarao et al., 2024) incurring inaccuracies. Finding ways to inject or update new and revised facts is crucial; Particularly for factuality-critical applications like medicine (Bao et al., 2023; Jeong et al., 2024), law (Colombo et al., 2024; Lai et al., 2024) or education (Team et al., 2024c; Wang et al., 2024b).
However, frequent full model retraining is prohibitively expensive. To address this, lifelong knowledge editing techniques (Hartvigsen et al., 2023; Meng et al., 2023; Wang et al., 2024a) have recently emerged. In controlled benchmark studies, they have shown success incorporating new facts along three key metrics: (1) generalization beyond merely memorizing question-answer pairs, (2) locality of edits to maintain pretraining knowledge, and (3) retention of all previously seen edits.
Unfortunately, existing benchmarks are small [^2] CounterFact (Meng et al., 2023) at 20K, ZsRE (Levy et al., 2017) at 1K, SelfCheckGPT (Manakul et al., 2023) at 600., synthetic, and precede the knowledge cutoff of modern LLMs. While more recent efforts Jang et al. (2022); Khodja et al. (2024) avoid synthetic edits by using wikidata, they remain small-scale (around 20K), and have yet to examine the difficulty of edits, limiting our understanding of really lifelong knowledge editing capabilities at practically-needed scales. LLMs are nowadays trained on trillions of text tokens (e.g. Llama-3 (et al., 2024a) on 15T), and studying lifelong factuality updates demands large volumes of recent real-world knowledge edits.
To better understand lifelong knowledge editing at a scale and scenario of practical importance, we introduce WikiBigEdit - a large-scale, lifelong benchmark (Prabhu et al., 2024; Ghosh et al., 2024) with a fully automated dataset extraction pipeline that continuously expands with new real-world factual edits to ensure future-proof factuality probing. At its inception, WikiBigEdit covers over 500K high-quality question-answer pairs (around 7 million tokens; magnitudes larger than standard edit benchmarks) spanning eight time-intervals over five months (February - July 2024) - reflecting a typical interval between subsequent version releases of large-scale LLMs [^3] E.g. Llama 1 (Touvron et al., 2023) to LLama 2 (et al., 2023), DeepSeek-V1 (et al., 2024b) to V2 (et al., 2024c) and DeepSeek-V3 (et al., 2024d). to better emulate the deployment time required to cover with factual updates.
Derived from periodic changes to Wikidata knowledge graphs (Jang et al., 2022; Khodja et al., 2024), WikiBigEdit covers a large range of factual edits and refinements. Moreover, WikiBigEdit introduces comprehensive evaluation axes going beyond standard knowledge editing performance measures, incorporating locality checks, multi-hop reasoning across incorporated edits as well as standard and complex generalization tests. We use WikiBigEdit to explore the properties of real-world knowledge edits and conduct an extensive study of knowledge editing techniques for WikiBigEdit-scale updates, focusing on their ability to integrate numerous sequential edits while maintaining previous knowledge and factual accuracy beyond simple memorization. Additionally, we compare these methods against other generic modification techniques such as retrieval augmentation (Vu et al., 2023; Prabhu et al., 2023a; Gui et al., 2024) and continual finetuning (Ibrahim et al., 2024; Roth et al., 2024). Our study reveals limited transfer of knowledge editing techniques to practically scalable lifelong updates and highlights that even when evaluated around knowledge editing desiderata, standard approaches such as retrieval augmentation or continual finetuning with model merging can perform better.
@@
  • We present WikiBigEdit: a large-scale benchmark of 500k high-quality question-answer pairs spanning eight timesteps over six months, derived from real-world factual changes in Wikidata to both analyze and simulate real-world knowledge edit tasks at a practically relevant scale.
  • A fully automated extraction pipeline which continuously extracts new, suitable factual edits from Wikidata knowledge graphs, enabling the benchmark to evolve with new knowledge updates continuously.
  • Using WikiBigEdit, we thoroughly analyze the capability of existing lifelong knowledge editing methods to conduct lifelong edits at scale; contrasted against retrieval augmentation and continual finetuning to understand limits in relation to other established approaches.
@@ 2 Related Works
NQ 307K 2016 Google Search queries Open factual QA
@@
EX-FEVER 60K 2023 Hyperlinked Wikipedia Mhop fact verif.
ELI5 270K 2019 Reddit (ELI5 subreddit) Long-form QA
@@
DatedData 200K Various Varied web sources Temporal QA
StreamingQA 150K 2007-2020 WMT news articles Event-based QA
@@
CLARK 1.4K 2021-2024 Wikipedia Knowledge QA
@@
TemporalWiki 7K 2021 Wikipedia, Wikidata Temporal QA
@@
WikiBigEdit 502K 2024 Wikidata Fact-based QA
@@
(Lifelong) Knowledge Editing Benchmarks. Traditional QA datasets like NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and MS MARCO (Bajaj et al., 2018) target open-domain factual QA, HotpotQA (Yang et al., 2018) and EX-FEVER (Ma et al., 2024) assess multi-hop reasoning. Fan et al. (2019); Li et al. (2024); Cheng et al. (2024); Mallen et al. (2023) focus on long-form explanatory questions and time-sensitive QA. StreamingQA (Liška et al., 2022), ArchivalQA (Wang et al., 2021) or HelloFresh (Franzmeyer et al., 2024) address evolving or time-sensitive, event-driven data, while TiC-LM (Li et al., 2025) explores continual pretraining of LLMs. Knowledge editing benchmarks assess the ability to integrate new facts without retraining. ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023) focus on synthetic updates, while AToKE (Yin et al., 2024) and ChronoEdit (Ge et al., 2024) evaluate temporal consistency but remain limited in scale and scope. ELKEN (Peng et al., 2024) and EVEDIT (Liu et al., 2024b) generate edits based on real-world events. TemporalWiki (Jang et al., 2022) and WikiFactDiff [^4] While WikiFactDiff contains over 450K triple–text pairs, only a 20K subset ($\text{WFD}_{\text{repl}}$) is commonly used for editing due to its defined evaluation protocol. (Khodja et al., 2024) are more closely aligned with WikiBigEdit. Both evaluate temporal update consistency using consecutive Wikidata snapshots. WikiBigEdit provides over a magnitude more samples (Table 2), future-proofing and a comprehensive evaluation along all editing desiderata.
@@
Knowledge Editing. Approaches for integrating new knowledge into LLMs include local modification methods such as ROME (Meng et al., 2023) and MEMIT (Meng et al., 2022), global optimization techniques like MEND (Mitchell et al., 2021), and external memorization strategies such as SERAC (Mitchell et al., 2022), GRACE (Hartvigsen et al., 2023), RECIPE (Chen et al., 2024), and WISE (Wang et al., 2024a). While these techniques support efficient model updates, they often suffer from catastrophic forgetting and knowledge degradation when applied to large-scale or sequential edits (Hsueh et al., 2024; Gupta et al., 2024b; Yang et al., 2024b). Our benchmark offers a practical framework to examine these limitations under realistic, evolving knowledge conditions.
Retrieval Augmented Generation (RAG). By combining a retrieval module with a generative language model, this class of methods addresses issues such as outdated knowledge and untraceable generations (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020). RAG approaches have shown strong performance on knowledge-intensive tasks, even at smaller model scales (Borgeaud et al., 2022; Izacard et al., 2022), and enable attribution by linking generated outputs to retrieved sources (Rashkin et al., 2022; Honovich et al., 2022). Moreover, RAG supports dynamic knowledge updates, mitigating context-memory conflicts when parametric information becomes stale (Vu et al., 2023). However, RAG faces limitations in multi-hop reasoning and compositional retrieval. More advanced variants - such as HippoRAG (Gutiérrez et al., 2025) or agent-based RAG systems - aim to address these limitations by improving reasoning over retrieved information.
@@
Continual Finetuning and Model Merging. Continual fine-tuning has emerged as a reliable approach for general large-scale model updates (Garg et al., 2024; Prabhu et al., 2023b; Ibrahim et al., 2024). Roth et al. (2024) show that parameter-efficient finetuning via e.g. LoRA (Hu et al., 2021) or GaLore (Zhao et al., 2024), can be very effective, leveraging regularization provided by limiting trainable parameter counts (Thede et al., 2024). Parameter-selective techniques, including BitFit (Zaken et al., 2022) and LNFit (Min et al., 2023), further reduce computational overhead by targeting specific parameter subsets. Additionally, regularization strategies from continual learning, such as EWC (Kirkpatrick et al., 2017) and SI (Zenke et al., 2017), have shown notable success in mitigating forgetting during the continual adaptation of foundation models. Model merging (Wortsman et al., 2022; Ilharco et al., 2022; Ramé et al., 2024) has shown promising gains coupled with finetuning for continual tasks (Stojanovski et al., 2022; Marouf et al., 2024; Kozal et al., 2024; Marczak et al., 2024; Dziadzio et al., 2024). We investigate and contrast these approaches against knowledge editing at scale.
@@ 3 WikiBigEdit
WikiBigEdit addresses three key aspects: (1) it extracts a large number of realistic knowledge edits, (2) incorporates extensive evaluations to assess generalization, reasoning, and locality, and (3) operates fully automatically. The complete pipeline is shown in Figure 1.
Problem Definition.
The task of large-scale, lifelong knowledge editing requires updating a pre-trained LLM $f^{0}$ with a stream of factual updates $u_{1},u_{2},\ldots,u_{T}]$, where each update $u_{t}=(q_{t},a_{t})$ consists of a question $q_{t}$ and its factual answer $a_{t}$. These updates are grouped into $B$ sequential batches $U_{1},U_{2},\ldots,U_{B}]$, where each batch can encompass anything from a single edit to multiple. For each batch update $b$ (also denoted as timestep in this work), the model $f^{b-1}$, trained on updates from prior batches $U_{<b}$, is further updated with the current batch $U_{b}$ to produce $f^{b}$. The model must meet three key objectives: (1) integrate updates from $U_{b}$, (2) retain knowledge from $U_{<b}$, and (3) ensure that the semantic understanding required to correctly solve previously seen edits generalizes beyond simple memorization of exact question-answer pairings.
@@ 3.1 Lifelong Automated Extraction of Factual Edits
As a proxy for world knowledge, we leverage Wikidata knowledge graphs, following recent works (Jang et al., 2022; Khodja et al., 2024). By performing pairwise comparisons of snapshots taken at different points in time, we identify changes that approximate real-world knowledge evolution. We thus automatically generate WikiBigEdit using our proposed seven-step pipeline, as follows:
@@
(5) Generation of Question-Answer Edit Pairs. For $\mathcal{S}_{\text{changed}}$, $\mathcal{S}_{\text{locality}}$, and $\mathcal{S}_{\text{mhop}}$, we generate corresponding question-answer pairs using GPT-4o-mini (Brown et al., 2020) (GPT-3.5 for update and locality) through few-shot prompting (for all prompt templates, please see the supplementary). For $\mathcal{S}_{\text{mhop}}$, we generate questions that omit the middle entity of the two-hop factual quintuple. Using this, we generate the corresponding question-answer sets $\mathcal{S}^{\text{QA}}_{\text{changed}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$. $\mathcal{S}^{\text{QA}}_{\text{locality}}$ and $\mathcal{S}^{\text{QA}}_{\text{mhop}}$ are used for evaluation, while $\mathcal{S}^{\text{QA}}_{\text{changed}}$ constitutes the respective fact-based training data.
@@ 3.2 Analysis and Exploration
In this section, we first summarize various benchmark characteristics before leveraging the first instantiation of our large-scale WikiBigEdit to study the structure and properties of real-world knowledge edits derived from Wikidata knowledge graphs. Table 2 provides an overview of the eight time-intervals that constitute WikiBigEdit. The factual updates are extracted on a roughly bi-monthly frequency from Wikidata snapshots spanning February to July 2024; with the availability of Wikidata snapshots determining size and extend of factual updates associated with each timestep.
To ensure high data quality, we filter out approximately 76% of initial changes, removing cyclic or overly long entities (11%), non-entity-to-entity relations (50%), and nondeterministic triplets (15%). The filtered updates range from 27k to 69k per step, with monthly factual changes exceeding 100k. This variability reflects the dynamic nature of knowledge updates and practical constraints when aiming to ensure that models remain up-to-date. Overall, this version contains 502,382 question-answer pairs across 6.9M tokens.
Impact of Language Model Choice on QA Generation The quality of generated question-answer pairs in WikiBigEdit depends in part on the underlying language model used during generation. Before we conduct further experiments, we begin by examining the sensitivity of the generation process to different LLMs by comparing outputs from GPT-3.5-turbo, GPT-4o-mini, and GPT-4o. For both the Update and Locality sets, which are directly derived from structured subject–relation–object triplets, model choice has negligible impact. Even the less capable GPT-3.5-turbo consistently produced well-formed, semantically correct questions. As shown in Table 3, the outputs across all models for such simple factual prompts are stylistically comparable and accurate. In contrast, the Multi-hop and Persona sets were more sensitive to model choice, as they require compositional reasoning and stylistic transformation. For instance, GPT-3.5-turbo often struggled to combine facts effectively in multi-hop settings, producing incomplete or ambiguous questions. GPT-4o-mini and GPT-4o, in comparison, reliably produced complete and coherent outputs. These observations confirm our selection for GPT-4o-mini to generate Multi-hop and Persona sets due to its strong balance between generation quality and computational efficiency.
Update Triplet Lea County Regional Airport - state of use - in use
GPT-3.5 What is the current state of use for Lea County Regional Airport?
GPT-4o-mini What is the current state of use of Lea County Regional Airport?
GPT-4o What is the current state of use of Lea County Regional Airport?
Mhop Tuple 2004 VS75 - discoverer - Marc Buie - gender - male
GPT-3.5 Who is the discoverer or inventor of 2004 VS75?
GPT-4o-mini What is the sex or gender of the discoverer of 2004 VS75?
GPT-4o What is the gender of the discoverer or inventor of 2004 VS75?
Table 3: Comparison of generated questions across different language models for a simple factual update (top) and a multi-hop reasoning example (bottom). While all models perform well on the single-hop case, GPT-3.5 struggles with compositional reasoning, often producing incomplete multi-hop questions. GPT-4o-mini and GPT-4o generate more complete and semantically accurate queries.
@@
Specificity of factual updates. We assess the specificity of knowledge edits by distinguishing well-known from highly specialized facts, using two measures: (1) entity frequency from the Infinigram API (Liu et al., 2024a), and (2) Wikipedia view counts (Jan 2023–June 2024) (Mallen et al., 2023). As shown in Figure 2 (d–g), WikiBigEdit includes both general (e.g., Which continent is Krasnodar a part of?) and highly specific questions (e.g., Who was the doctoral advisor of Omar Khayyám?). To evaluate how specificity affects performance, we test five LLMs (Llama-2-7b, Llama-3-8b, Mistral-7b, Gemma-7b, xGen-7b) and correlate their accuracy. As shown in (h), lower entity frequency correlates with reduced performance, consistent with Udandarao et al. (2024); Cheng et al. (2024), while view counts (i) show less consistent trends. Specificity appears primarily driven by the object rather than the subject. See supplementary for details.